OpenStudy (anonymous):

Find the coordinates of the midpoint of a segment with the endpoints (16,5) & (28, -13).

OpenStudy (ddcamp):

Midpoint formula: \[M = \left( \frac{ x₂-x₁ }{ 2 }, \frac{ y₂-y₁ }{ 2 } \right)\]

OpenStudy (anonymous):

what did you get when you solved? i didnt get it

OpenStudy (ddcamp):

So you have your 2 points, (x₁,y₁) and (x₂,y₂), or in this problem (16,5) and (28,-13) x₁ = 16 x₂ = 28 y₁ = 5 y₂ = -13 Just plug those values into the midpoint formula and you'll get the point M

OpenStudy (anonymous):

but i got 6,9 but thats not in the answer choices

OpenStudy (ddcamp):

Is the 9 positive or negative?

OpenStudy (anonymous):

-9

OpenStudy (ddcamp):

It should be (6,-9), at least, that's what I got.

OpenStudy (anonymous):

i got that too but its not one of the choices

OpenStudy (ddcamp):

<facepalm> I copied the formula wrong, it should be adding, not subtracting.

OpenStudy (anonymous):

okay so what did you get?

OpenStudy (ddcamp):

(22, -4)